find the equation of a line through (6,-8), perpendicular to 9y=x-18
Answer by lwsshak3(11628)

find the equation of a line through (6,-8), perpendicular to 9y=x-18
**
Standard form of equation for a straight line{ y=mx+b, m=slope, b=y-intercept.
9y=x-18
y=x/9-2
m=1/9
..
slope of perpendicular line =-9 (negative reciprocal of 1/9)
Equation:
y=-9x+b
solve for b using coordinates of given point (6,-8) thru which line passes.
-8=-9*6+b
b=54-8=46
Equation of perpendicular line:
y=-9x+46